### v3

@@ Abstract
The weather forecasting system is important for science and society, and significant achievements have been made in applying artificial intelligence (AI) to medium-range weather forecasting. However, existing AI-based weather forecasting models rely on analysis or reanalysis products from traditional numerical weather prediction (NWP) systems as initial conditions for making predictions. Initial states are typically generated by traditional data assimilation components, which are computational expensive and time-consuming. Here we present an AI-based data assimilation model, i.e., Adas, for global weather variables. By introducing the confidence matrix, Adas employs gated convolution to handle sparse observations and gated cross-attention for capturing the interactions between the background and observations. Further, we combine Adas with the advanced AI-based forecasting model (i.e., FengWu) to construct the first end-to-end AI-based global weather forecasting system: FengWu-Adas. We demonstrate that Adas can assimilate global observations to produce high-quality analysis, enabling the system operate stably for long term. Moreover, we are the first to apply the methods to real-world scenarios, which is more challenging and has considerable practical application potential. We have also achieved the forecasts based on the analyses generated by AI with a skillful forecast lead time exceeding that of the IFS for the first time.
@@
Driven by the advancements and maturity of AI, particularly deep learning techniques, scientific intelligence has been rapidly evolving with the aim of leveraging AI to promote scientific research and discovery. Within the field of atmospheric science, AI has achieved remarkable achievements in various areas, such as post-processing and bias correction [1, 2, 3, 4], downscaling [5, 6], precipitation nowcasting [7, 8, 9], climate forecasting [10], and medium-range weather forecasting [11, 12, 13, 14, 15]. Some AI-based models have demonstrated highly competitive forecasts compared to the deterministic forecasts of the state-of-the-art NWP system, i.e., the Integrated Forecasting System (IFS) from the European Centre for Medium-Range Weather Forecasts (ECMWF). These models are usually trained on reanalysis dataset (e.g., ERA5 [16] from ECMWF) and allow much lower computational costs and easier deployment for operational forecasting. Despite drawbacks such as forecast smoothness and bias drift [17], AI approaches have shown the immense potential of data-driven modeling in weather prediction, offering a new paradigm for meteorological forecasting.
@@
However, the AI-based weather forecasting models still require analysis products generated through the process of data assimilation in the traditional NWP system for making predictions (Fig. 1). Specifically, data assimilation aims to obtain the best estimate of the true state of the Earth system (known as the analysis) and provide an accurate initial state for weather prediction, thereby improving the forecast performance. With periodic corrections applied to the forecasts, data assimilation allows for a more stable and accurate long-term prediction, which is a critical component in global weather forecasting system. This process relies heavily on observations as they provide crucial information that closely represents the true state of the atmosphere. For example, the earliest initial conditions were obtained by interpolating observations onto the grid points of the grid space [18]. Modern data assimilation techniques are usually achieved by integrating observations with the short-range forecasts (i.e., the background), primarily including two categories: Kalman filters and variational methods [19]. The main motivation for introducing the background, or the first guess, lies in the incomplete nature of observations. The sparsity of observations makes it difficult to obtain a global estimate, resulting in low-quality interpolation analysis. Besides, not all state variables can be observed directly.
@@
Currently, AI methods have primarily been applied to specific processes within data assimilation to address the limitations of traditional data assimilation techniques [20, 21]. For instance, implicit neural representations (INRs) [22] and various autoencoders (AEs) [23, 24, 25] have been employed as reduced-order-models (ROMs) to tackle high-dimensional challenges. These methods replace classical linear models such as Proper Orthogonal Decomposition (POD) [26, 27] in latent assimilation, providing an efficient framework for the representation and reconstruction of state variables in the latent space. Neural networks have also been utilized to derive tangent-linear and adjoint models in 4D variational (4D-Var) data assimilation [28], provide localization functions for ensemble Kalman filters (EnKF) [29] and estimate error covariance matrix [30, 31]. Moreover, there exist strong mathematical similarities between machine learning and data assimilation, particularly the variational data assimilation [32], which enables the optimization of the cost function in variational data assimilation using auto-differentiation [24, 33]. However, these studies have not involved altering the algorithms of data assimilation fundamentally. While traditional data assimilation algorithms are underpinned by rigorous mathematical theory and physical priors, they inevitably rely on certain assumptions as prerequisites. In practical operational applications, they often face constraints due to high computational costs, necessitating various approximations for solution [34]. In situations where supervised information is available, neural networks are expected to capture the correlations among data during the assimilation automatically and provide a new kind of algorithms [35]. Efforts have been made in some simple scenarios [36, 37, 38], but they simply feed the data into networks without any specialized design. And so far, there have been no attempts in data assimilation for real observational data in multivariate global weather scenarios.
@@
In this study, we present Adas, a novel data assimilation model for global weather variables, which can assimilate sparse observations with different qualities and provide initial conditions for weather forecasting. Inspired by the gating mask in gated convolution [39] and the error covariance matrix in traditional data assimilation techniques, we introduce the confidence matrix to characterize the availability and quality of observations during the data assimilation. With the guidance of confidence matrix, Adas employs the gated convolution module to handle sparse observations and the gated cross-attention module for capturing the interactions between the background and observations efficiently. The prediction of the advanced AI-based weather forecasting model, FengWu [14], is used to generate the background for data assimilation. In return, the analysis of Adas is then used as the initial state of FengWu for making predictions at the next time step, thus forming the first end-to-end AI-based global weather forecasting system: FengWu-Adas. Fig. 1 shows the progression of global weather forecasting system and the main motivation of this research. We demonstrate that our system can operate stably for long term, and our method is robust and flexible to be combined with any state-of-the-art weather forecasting model without retraining or fine-tuning. The simulation experiments have proven the superior performance of FengWu-Adas in capturing the distribution patterns of the background error and producing high-quality analysis. And we further apply FengWu-Adas to real-world scenarios, which is more challenging and has considerable practical application potential. We have also achieved the forecasts based on the analyses generated by AI with a skillful forecast lead time of more than 8.75 days, which exceeds that of the IFS for the first time.
@@ Results
The framework of FengWu-Adas is shown schematically in Fig. 2a. For the given initial state denoted as $\dot{x}_{t}$ at time $t$, FengWu can make multi-step predictions $\hat{x}_{t+\Delta t}$, $\hat{x}_{t+2\Delta t}$, …, $\hat{x}_{t+k\Delta t}$ at time $t+\Delta t$, $t+2\Delta t$, …, $t+k\Delta t$ in an auto-regressive manner, where $k$ denotes the time index and $\Delta t=6h$ is the temporal spacing of single-step prediction. Considering that the accuracy of the background will directly affect the quality of the analysis, the single-step prediction of FengWu is used as the background $x_{t+\Delta t}^{b}$ for data assimilation. After obtaining the observations $y_{t+\Delta t}$ at the corresponding time, confidence levels ranging from 0 to 1 can be assigned to each datum based on the availability and quality of the observation, thereby generating the confidence matrix. The confidence matrix $m$, serving as supplementary information, is sent into Adas along with observations for assimilation. Subsequently, the analysis $x_{t+\Delta t}^{a}$ produced by Adas will be the initial state for the next prediction, i.e., the input of FengWu at next time and so on, in a cyclic manner. The multi-step forecasts are predicted through auto-regression based on the analysis.
@@
Adas utilizes a dual encoder composed of gated interaction blocks to extract features of the background and sparse observations separately and capture the interactions between them (Fig. 2b). All the inputs with a shape of $H_{0}\times W_{0}\times C_{0}$ are converted into a unified latent shape of $D\times H\times W\times C$ through the patch embedding module so that the network can capture their spatial relationships in the 3D mesh. $H_{0}$ and $W_{0}$ represent the number of grid points in the longitudinal and latitudinal directions, respectively, and the channel $C_{0}$ represents the number of variables. The 3D upper-air variables and 2D surface variables are embedded into the latent space independently and then concatenated together, where the information at each grid point is represented by a $C$-dimensional vector. $H$ and $W$ correspond to the number of grid points in the longitudinal and latitudinal directions after embedding, and $D$ represents the number of levels in the vertical direction. By performing down-sampling and up-sampling through the patch merging and patch expanding module (Fig. 2d), the encoder can capture multi-scale meteorological features. The dual gated interaction block extracts features from the background and observations with standard convolution and gated convolution, along with efficient information interactions through the gated cross-attention module (Fig. 2c). Both the gated convolution and gated cross-attention are guided by the confidence matrix, which is updated together with the features to represent the availability and quality of observations (details in Methods). After the feature fusion block implemented by a convolutional layer, the features are sent into a series of Transformer [40, 41] blocks and then recovered to the original size. The convolutional operation can bring inductive bias of locality, and the attention mechanism enables the model to capture long-range dependencies. The attention block utilizes the shifted-window mechaism [42] adjusted according to geographical rules to reduce the computational cost and ensure continuity in the latitudinal direction, and $N=8$ in our implementation.
Superior performance of FengWu-Adas in simulation experiments
For testing the performance of FengWu-Adas, the Observing System Simulation Experiments (OSSEs) are performed by utilizing the ERA5 dataset firstly. The experiment proceeds with alternating weather forecasting and data assimilation steps: the single-step forecast provides the background for data assimilation, and the analysis is used as the initial state for the next forecasting step. ERA5 is regarded as the ground truth, and the simulated observations are obtained by adding a mask to the ground truth. The mask is used as the simulated confidence matrix for data assimilation. Therefore, the confidence matrix in simulation experiments is binary, with 1 indicating that the observation in corresponding grid point is available and accurate, and 0 indicating unavailable observation. The confidence matrix in real scenarios will be more complex and the confidence value is produced based on the quality of the observations. To improve the robustness of observation locations, the simulated masks for training are generated randomly with a fixed mask ratio at each time step. However, most observation locations are generally fixed in real-world scenarios, and the same strategy for testing not only deviates from reality but also poses potential issues in the cyclic experiments. Specifically, a random mask implies that any position has the probability of being observed. When the number of steps in the cyclic experiments is sufficiently large, the observation positions would approximately cover the whole world, which is clearly unrealistic. To address this concern, the masks for testing are fixed even though we find that this has almost no impact on performance in our experiments (see the comparison between Fig. 3 and Extended Data Fig. 7).
Here, we conduct three sets of experiments with different observation ratios, i.e., 90%, 99% and 99.9%. In other words, the ratios of available observations to the total number of grid points are 10%, 1% and 0.1%. The system is initialized using the ERA5 data at 2018-01-01 00:00UTC, which serves as the initial state for the first-step forecasting. Once the system is initialized, ERA5 data is invisible throughout the whole year, except for generating simulated observations. Fig. 3 illustrates the variations in RMSE and Bias (defined in Methods) of the analyses for main variables over the year. Both RMSE and Bias are computed against ERA5. The results with different observation ratios are represented by different colors, as marked in the bottom. To enhance clarity, the original data is presented with reduced opacity, while the solid lines indicate the smoothed values after the exponential moving average (EMA) with a smoothing factor of 90%. The RMSE of the analyses for all variables with all observation ratios is maintained at a low level for the whole year, showing the long-term stability of the system and the ability to generate accurate analysis based on Adas. The dashed lines are the RMSE of 6-hour forecast of the IFS, which represent fairly advanced error levels. Even when the observation ratio is only 0.1% (corresponding to an absolute number of observations of less than 0.1 million), the RMSE for all variables is lower than the dashed lines at almost all times. As the observation ratio grows, the RMSE of the analysis decreases gradually, and the Bias remains at around 0 more stably, revealing the system’s potential to eliminate bias drift of AI-based forecasting models and maintain conservation. The average RMSE over the whole year for all variables is shown in Extended Data Fig. 8.
The visualizations for z500 during the data assimilation, along with the error distributions and analysis increments, are presented in Fig. 4. The visualization date-time is randomly selected at 2018-01-26 06:00UTC, where the background is obtained from FengWu by performing single-step forecasting of the analysis at 2018-01-26 00:00UTC. The first row displays the visualizations of ERA5 (i.e., the ground truth, Fig. 4a) and the background (Fig. 4b), and the error distributions of the background (i.e., the background minus ERA5, Fig. 4c). The other rows show the visualizations of random simulated observations (the first column), the analysis increments (i.e., the analysis minus the background, the second column), and the error distributions of the analysis after data assimilation (i.e., the analysis minus ERA5, the last column) with different observation ratios. All the analysis increments exhibit a distribution similar to the background errors, indicating that Adas captures the error distribution of the background effectively. And all the errors of the analysis are reduced to varying degrees after data assimilation. As expected, the analysis increment is closer to the background error and the analysis error is smaller with the increase of observation information.
In order to further demonstrate the stability of the system, a completely random initial state is used to start the system. Extended Data Fig. 9 shows the corresponding RMSE, and the subgraph provides a clearer display of the first ten days. The y-axis in these graphs uses a coordinate axis with non equidistant logarithmic form to ensure visual perception. The first-step forecast after random initialization will produce a terrible performance, but all the errors converge quickly to a stable state with the same levels of metrics as that of starting the system with ERA5 in a few days. And the convergence speed accelerates as the number of observations increases, which is reasonable. This indicates that Adas can perform data assimilation on the background with any quality, ensuring that the system can correct back to the stable state in a short period of time without affecting subsequent operations even if it encounters a malfunction that leads to complete derailment of the forecast.
As an end-to-end global weather forecasting system that can operate independently based on observational data, FengWu-Adas can generate multi-step forecasts on its own analysis (Fig. 2a). The average RMSE and ACC skill for weather forecasting of FengWu-Adas over ten days is shown in Extended Data Fig. 10. It should be emphasized that FengWu is trained on ERA5 data [14] and is not fine-tuned to our task. Therefore, our method is flexible to be combined with any state-of-the-art weather forecasting model. Actually, our pre-trained model can assimilate observations on arbitrary background without retraining or fine-tuning, as shown in Extended Data Fig. 9. The universal capability allows Adas to be directly integrated with other weather forecasting models as a plug-and-play component, and the forecasting skill of the system will improve as the forecasting model develops.

### v4

@@ Abstract
The weather forecasting system is important for science and society, and significant achievements have been made in applying artificial intelligence (AI) to medium-range weather forecasting. However, existing AI-based weather forecasting models rely on analysis or reanalysis products from traditional numerical weather prediction (NWP) systems as initial conditions for making predictions. The initial states are typically generated by traditional data assimilation components, which are computationally expensive and time-consuming. Here, by cyclic training to model the steady-state background error covariance and introducing the confidence matrix to characterize the quality of observations, we present an AI-based data assimilation model, i.e., Adas, for global weather variables. Further, we combine Adas with the advanced AI-based forecasting model (i.e., FengWu) to construct an end-to-end AI-based global weather forecasting system: FengWu-Adas. We demonstrate that Adas can assimilate global conventional observations to produce high-quality analysis, enabling the system to operate stably for long term. Moreover, the system can generate accurate end-to-end weather forecasts with comparable skill to those of the IFS, demonstrating the promising potential of data-driven approaches.
1 Introduction
@@
Driven by the advancements and maturity of AI, particularly deep learning techniques, scientific intelligence has been rapidly evolving with the aim of leveraging AI to promote scientific research and discovery. Within the field of atmospheric science, AI has achieved remarkable achievements in various areas, such as post-processing and bias correction (yang2023improving; mouatadid2023adaptive; agrawal2023machine; singh2022short), downscaling (harder2022generating; chen2022rainnet), precipitation nowcasting (ravuri2021skilful; gao2022earthformer; zhang2023skilful), climate forecasting (ling2022multi), and medium-range weather forecasting (pathak2022fourcastnet; bi2023accurate; lam2023learning; chen2023fengwu; chen2023fuxi). Some AI-based models have demonstrated highly competitive forecasts compared to the deterministic forecasts of the state-of-the-art NWP system, i.e., the Integrated Forecasting System (IFS) from the European Centre for Medium-Range Weather Forecasts (ECMWF). These models are usually trained on reanalysis dataset (e.g., ERA5 (hersbach2020era5) from ECMWF) and allow much lower computational costs and easier deployment for operational forecasting. Despite drawbacks such as forecast smoothness, bias drift and lack of physical consistency (ben2023rise; bonavita2024some), AI approaches have shown the immense potential of data-driven modeling in weather prediction, offering a new paradigm for meteorological forecasting.
@@
However, the AI-based weather forecasting models still require analysis products generated through the process of data assimilation in the traditional NWP system for making predictions (Figure 1). Specifically, data assimilation aims to obtain the best estimate of the true state of the Earth system (known as the analysis) and provide an accurate initial state for weather prediction, thus improving the forecast performance. In the operational weather forecasting system, data assimilation is a critical component that ensures the long-term stable operation of the system by periodically correcting the forecast. Beyond the predictability limit, which is believed to be about 15 days, the forecast results without data assimilation will become completely unreliable. This process is heavily dependent on observations, as they provide crucial information that closely represents the true state of the atmosphere. For example, the earliest initial conditions were obtained by interpolating observations onto the grid points of the grid space (richardson1922weather). Modern data assimilation techniques are usually achieved by integrating observations with short-range forecasts (i.e., the background), primarily including two categories: Kalman filters and variational methods (rabier2003variational). The main motivation for introducing the background, or the first guess, lies in the incomplete nature of observations. The sparsity and irregularity of observations make it difficult to obtain a global estimate, resulting in low-quality interpolation analysis. Besides, not all instruments can observe geophysical variables directly.
Figure 1: The progression of global weather forecasting system. Traditional NWP systems consist of physical weather forecasting model and data assimilation. The breakthrough of AI-based medium-range weather forecasting models occurred in 2022-2023 with highly competitive performance in terms of accuracy, but they still rely on the NWP systems for making predictions. Our work is dedicated to exploring the possibility of an end-to-end global weather forecasting system which is driven purely by AI.
@@
Currently, AI methods have been applied primarily to specific processes within data assimilation to address the limitations of traditional data assimilation techniques (buizza2022data; cheng2023machine). For example, implicit neural representations (INRs) (li2023latent) and various autoencoders (AEs) (peyron2021latent; melinc2023neural; amendola2021data) have been employed as reduced-order models (ROMs) to address high-dimensional challenges. These methods replace classical linear models such as proper orthogonal decomposition (POD) (cheng2023generalised; pawar2022equation) in latent assimilation, providing an efficient framework for the representation and reconstruction of geophysical variables in latent space. Neural networks have also been utilized to derive tangent-linear and adjoint models in 4D variational (4DVar) data assimilation (hatfield2021building), provide localization functions for ensemble Kalman filters (EnKF) (wang2023convolutional) and estimate error covariance matrix (cheng2022observation; penny2022integrating). Moreover, there exist strong mathematical similarities between machine learning and data assimilation, particularly the variational data assimilation (geer2021learning), which enables the optimization of the cost function in variational data assimilation using auto-differentiation (melinc2023neural; xiao2023fengwu). However, these studies have not involved fundamentally altering the algorithms of data assimilation. While traditional data assimilation algorithms are underpinned by rigorous mathematical theory and physical models, they inevitably rely on certain assumptions as prerequisites. In practical operational applications, they often face constraints due to high computational costs, which requires various approximations for solution (carrassi2018data). In situations where supervised information is available, neural networks are expected to automatically capture the correlations among data during the assimilation and provide a new alternative algorithm (fablet2021learning). Efforts have been made in idealized data assimilation experiments (rozet2023score; huang2024diffda; chen2024fnp), and further research focuses on data-driven end-to-end weather forecasting systems, broadly on systems that can initialize forecasts without needing traditional physically-based analysis (vaughan2024aardvark; xu2024fuxi; mcnally2024data). Some follow the paradigm of traditional NWP systems and specially design data assimilation models to work together with AI-based weather forecasting models in a forecast-assimilation cycle to form a system (xu2024fuxi). Others implicitly integrate the process of data assimilation and weather forecasting, using a single model to make predictions directly from observations (vaughan2024aardvark; mcnally2024data).
@@
In this study, we present Adas, a novel data assimilation model for global weather variables, which can assimilate sparse conventional observations with different qualities and provide initial conditions for weather forecasting. The key of data assimilation lies in adjusting the weights of the background and observations based on error covariance, and traditional data assimilation techniques primarily derive optimal analysis under the Bayesian framework. In our method, we model the steady-state (i.e., optimal dynamic balance between forecasting error growth and assimilation error reduction) background error covariance by cyclic training and introduce the confidence matrix to characterize the availability and quality of observations during the data assimilation. With the guidance of confidence matrix, Adas employs the gated convolution module to handle sparse observations and the gated cross-attention module for capturing the interactions between the background and observations efficiently. The prediction of the advanced AI-based weather forecasting model, FengWu (chen2023fengwu), is used to generate the background for data assimilation. In return, the analysis of Adas is then used as the initial state of FengWu for making predictions at the next time step, thus forming an end-to-end AI-based global weather forecasting system: FengWu-Adas. We demonstrate that our system can operate stably for long term, and our method is robust and flexible to be combined with any state-of-the-art weather forecasting model without retraining or fine-tuning. The ideal experiments have demonstrated the superior properties of FengWu-Adas in producing accurate analysis. Further, we apply FengWu-Adas to real-world scenarios, conducting a comprehensive and detailed evaluation to the system and making comparison with the IFS. The results show that FengWu-Adas can generate high-quality analysis and accurate end-to-end global weather forecasts although it only uses conventional observations, demonstrating the promising potential of data-driven approaches.
@@ Results
The framework of FengWu-Adas is shown schematically in Figure 2a. For the given initial state denoted as $\dot{x}_{t}$ at time $t$, FengWu can make multi-step predictions $\hat{x}_{t+\Delta t}$, $\hat{x}_{t+2\Delta t}$, …, $\hat{x}_{t+k\Delta t}$ at time $t+\Delta t$, $t+2\Delta t$, …, $t+k\Delta t$ in an auto-regressive manner, where $k$ denotes the time index and $\Delta t=6h$ is the temporal spacing of single-step prediction. Considering that the accuracy of the background will directly affect the quality of the analysis, the single-step prediction of FengWu is used as the background $x_{t+\Delta t}^{b}$ for data assimilation. After obtaining the observations $y_{t+\Delta t}$ at the corresponding time, confidence levels ranging from 0 to 1 can be assigned to each value in observation field based on the availability and quality of the observation, thereby generating the confidence matrix. The confidence matrix $m$, serving as additional information for the uncertainty of observation error, is sent into Adas along with observations for assimilation. For Adas, the input dataset is the background and observations, while the target dataset is ERA5, and the learning objective is to integrate the background forecast and observations to obtain analysis close to ERA5. Subsequently, the analysis $x_{t+\Delta t}^{a}$ produced by Adas will be the initial state for the next prediction, i.e., the input of FengWu at next time and so on, in a cyclic manner. Different from the general single-step training with a fixed lead background, this cyclic training makes the model more robust to the background, because the background quality by starting the forecast with analysis at the previous time changes dynamically with the convergence of the model training. By continuously iterating to optimize the implicit modeling of background error covariance, the system can achieve a steady state that balances the error growth of forecasting and the error reduction of assimilation. This approach is not limited to the optimal analysis of single-step assimilation, instead seeking a dynamic equilibrium for the forecasting and assimilation of the entire system, whose superiority has been previously substantiated (yang2006data; kalnay20074). As an end-to-end global weather forecasting system, FengWu-Adas can also perform multi-step forecasts through auto-regression based on the analysis.
Figure 2: The framework of FengWu-Adas and network architecture of Adas. a, The framework and process of FengWu-Adas. The prediction of FengWu provides the background for data assimilation and the analysis of Adas serves as the initial state for weather forecasting. The parameters of FengWu are frozen during training and not further fine-tuned or involved in joint optimization. b, Overview of Adas’s network structure. The dual encoder composed of gated interaction blocks extract multi-scale features of the background and sparse observations separately and capture the interactions between them, and then the features are fused and decoded. c, Details of dual gated interaction block. The gated convolution and gated cross-attention modules are designed for sparse observations, which are both guided by the confidence matrix. d, Simple schematic diagram of patch merging and patch expanding. The patch merging performs the rearrangement operation and linear layer to achieve down-sampling and the patch expanding performs the opposite operation to achieve up-sampling.
@@
Adas utilizes a dual encoder composed of gated interaction blocks to extract features of the background and sparse observations separately and capture the interactions between them (Figure 2b). All the inputs with a shape of $H_{0}\times W_{0}\times C_{0}$ are converted into a unified latent shape of $D\times H\times W\times C$ through the patch embedding module so that the network can capture their spatial relationships in the 3D mesh. $H_{0}$ and $W_{0}$ represent the number of grid points in the longitudinal and latitudinal directions, respectively, and the channel $C_{0}$ represents the number of variables. The 3D upper-air variables and 2D surface variables are embedded into the latent space independently and then concatenated together, where the information at each grid point is represented by a $C$-dimensional vector. $H$ and $W$ correspond to the number of grid points in the longitudinal and latitudinal directions after embedding, and $D$ represents the number of levels in the vertical direction. By performing down-sampling and up-sampling through the patch merging and patch expanding module (Figure 2d), the encoder can capture multi-scale meteorological features. The dual gated interaction block extracts features from the background and observations with standard convolution and gated convolution, along with efficient information interactions through the gated cross-attention module (Figure 2c). Both the gated convolution and gated cross-attention are guided by the confidence matrix, which is updated together with the features to represent the availability and quality of observations (details in Methods). After the feature fusion block implemented by a convolutional layer, the features are sent into a series of Transformer (vaswani2017attention; dosovitskiy2020image) blocks and then recovered to the original size. The convolutional operation can bring inductive bias of locality, and the attention mechanism enables the model to capture long-range dependencies. The attention block utilizes the shifted-window mechanism (liu2021swin) adjusted according to geographical rules to reduce the computational cost and ensure continuity in the latitudinal direction.
Superiority of FengWu-Adas in ideal experiments
The ideal cyclic forecast-assimilation experiments are first conducted using the ERA5 dataset to verify the performance and characteristics of FengWu-Adas. The experiments proceed with alternating weather forecasting and data assimilation steps: the single-step forecast serves as the background for data assimilation, and the analysis provides the initial state for the next forecasting step. ERA5 is considered as the ground truth, and simulated observations are obtained by sampling it with binary mask, which is then used as the confidence matrix for data assimilation. For simplicity, all variables are sampled independently in a certain proportion. Simulated masks for training are generated randomly at each time step to enhance robustness with respect to the observation locations. During the testing period, however, the masks remain fixed to avoid spurious performance caused by variations in information coverage at different times. The system is initialized with a 5-day lead forecast as the background, and except for producing simulated observations, ERA5 data is invisible throughout the entire experiments. Figure 3a illustrates the RMSE (defined in Methods) variations of the analyses for $z500$ and $t850$ variables over the year, and the RMSE is calculated against ERA5. The results with different observation ratios are represented by different colors, as marked in the legend. For clarity, the original data is presented with reduced opacity, while solid lines indicate the smoothed values after exponential moving average (EMA). The RMSE of the analyses with all observation ratios rapidly decreases and remains at a low level for the whole year, showing the long-term stability of the system and its ability to generate accurate analysis based on Adas. It should be emphasized that FengWu is trained on ERA5 data (chen2023fengwu) and is not fine-tuned to our task. As the observation ratio (i.e., the amount of observation information) increases, the speed at which the analysis RMSE decreases becomes faster, and the analysis RMSE in the final steady state is also lower, which is consistent with our general understanding.
Figure 3: The performance and properties of FengWu-Adas in ideal experiments. The RMSE (lower is better) is calculated against ERA5. a, RMSE variations of the analyses for $z500$ and $t850$ variables with different observation ratios in a whole year. The forecast with a 5-day lead time is used as the initial background to start the system. As the system continuously executes the cyclic forecast-assimilation, the RMSE of the analysis decreases rapidly and converges to a lower level. b, The performance differences between cyclic training and single-step training. The cyclic training directly models the background error covariance at steady state, circumventing the issue of error accumulation present in single-step training during cyclic forecast-assimilation experiments and achieving lower steady-state errors. c, The performance when starting the system with completely random Gaussian noise as the initial background. The system can still quickly recover to normal steady-state levels, showing outstanding stability and generalization capability to assimilate observations on arbitrary background. d, The power spectra of the 10-day lead background and analysis for $z500$ variable. After data assimilation, Adas can significantly correct the spectral degradation problem of FengWu forecasts with long lead time at small and medium scales. e, The sensitivity of the system to the noises with varying magnitudes. The horizontal axes correspond to different random Gaussian noise proportions of 0, 0.1%, 0.5%, 1%, 2.5%, 5%, and 10%, respectively. It demonstrates the robustness and anti-interference capability of Adas in the presence of noise.
We compare the RMSE variations of the analyses under single-step training and cyclic training to demonstrate the effectiveness of the training method in Figure 3b. The cyclic training directly models the background error covariance at steady state, circumventing the issue of error accumulation present in single-step training during cyclic forecast-assimilation experiments and achieving lower steady-state errors. Moreover, as the model encounters a more diverse range of background samples during training, it exhibits enhanced robustness to backgrounds with varying qualities and results in reduced error fluctuations. With this favorable property, even when starting the system with completely random Gaussian noise as the initial background, it can converge rapidly to the same steady-state level within a few days, and the convergence speed accelerates as the number of observations increases (Figure 3c). This ensures that the system can correct back to the stable state in a short period of time without affecting subsequent operations even if it encounters a malfunction that leads to complete derailment of the forecast. The generalization capability allows Adas to assimilate observations on arbitrary background without retraining or fine-tuning, so it is flexible to be integrated with any state-of-the-art weather forecasting model as a plug-and-play component.
A major challenge of current AI-based forecasting models is the smoothness of their predictions, which often manifests itself as degraded spectra (kochkov2024neural; lang2024aifs). Preserving fine-scale information is also important for data assimilation, allowing for better representation of errors at small and medium scales, which then grow to affect larger scales. Figure 3d shows the power spectra of the 10-day lead background and analysis for $z500$ variable. After data assimilation, Adas can significantly correct the spectral degradation problem of FengWu forecasts with long lead times at small and medium scales. And as the amount of observation information increases, the spectra of analysis will be closer to that of ERA5. In addition, we examine the sensitivity of the system to the noises with varying magnitudes and the balance problem when assimilating only temperature observations. Figure 3e shows the changes in RMSE of $z500$ and $t850$ with random Gaussian noise proportions of 0, 0.1%, 0.5%, 1%, 2.5%, 5%, and 10%, respectively. The change in analysis error caused by observational error is much smaller than the observational error itself, demonstrating the robustness and anti-interference capability of Adas in the presence of noise. As the number of observations increases, the model’s sensitivity to observational error will decrease further.